OpenStudy (anonymous):

Please Help! An open box is formed by cutting a 5 inch square measured from each corner and folding up the sides. If the volume of the carton is then 40 in3, what was the length of a side of the original square of cardboard

OpenStudy (anonymous):

you see the box is made by cutting 5 inch square from each corner, so the height of the box will be 5 inches, now volume of box is given as 40 in3 so area of base of the box = volume/ height = (40/5) in2 = 8 in2 now that you have area, and its given that originally the cardboard was square, hence length of side of square = \[\sqrt{8} in\]\[2\sqrt{2} in\]

OpenStudy (anonymous):

ohh sorry that will be the length of a side of the box, the length of actual cardboard will be \[(2\sqrt{2} + 10) in\]the 10 being added to make up for the 5 in squares cut from corners.